# V. Mehrmann and E. MengiRobust Stability of a Quadratic Matrix Polynomial

### v2

@@ 1 Introduction
(2) $\begin{split}&M(\nu)\;=\;\>\psi_{1}(\nu)M_{1}+\dots+\psi_{\mathsf{m}}(\nu)M_{\mathsf{m}}\,,\\[3.00003pt] &C(\nu)\;\>=\;\>\zeta_{1}(\nu)C_{1}+\dots+\zeta_{\mathsf{c}}(\nu)C_{\mathsf{c}}\,,\\[3.00003pt] &K(\nu)\;\,=\;\>\kappa_{1}(\nu)K_{1}+\dots+\kappa_{\mathsf{k}}(\nu)K_{\mathsf{k}}\,,\end{split}$
for given (parameter-independent) matrices $M_{i},C_{j},K_{\ell}\in{\mathbb{C}}^{\mathsf{n}\times\mathsf{n}}$, and real-analytic functions $\psi_{i},\zeta_{j},\kappa_{\ell}:\Omega\rightarrow{\mathbb{R}}$ for $i=1,\dots,\mathsf{m}$, $j=1,\dots,\mathsf{c}$, $\>\ell=1,\dots,\mathsf{k}$. In a context where the dependence of $P(\lambda;\nu)$ on $\nu$ is not important, we may leave out this explicit dependence on $\nu$, and simply write
@@ 2.1 Computation of the $\epsilon$-pseudospectral abscissa for small-scale problems
    ${\mathcal{M}}_{j}(x)\;:=\;\left\{\begin{array}[]{cl}\left[\begin{array}[]{cc}-\epsilon p^{2}_{w}(x)I&\;[P(x)]^{\ast}\\[2.5pt] P(x)&-\epsilon I\end{array}\right]&\quad\text{if }j=0\;,\\[13.00005pt] \left[\begin{array}[]{cc}0&\;-[P^{\prime}(x)]^{\ast}\\[2.5pt] P^{\prime}(x)&\;0\end{array}\right]&\quad\text{if }j=1\;,\\[16.00008pt] \left[\begin{array}[]{cc}\epsilon(2w_{m}^{2}x^{2}+w_{c}^{2})I&\;M^{\ast}\\[2.5pt] M&0\end{array}\right]&\quad\text{if }j=2\;,\\[18.99995pt] 0&\quad\text{if }j=3\;,\\[13.00005pt] \left[\begin{array}[]{cc}-\epsilon w_{m}^{2}I&\;0\\ 0&\;0\end{array}\right]&\quad\text{if }j=4\end{array}\right.$
@@
${\mathcal{N}}_{j}(y)\;:=\;\left\{\begin{array}[]{cl}\left[\begin{array}[]{cc}-\epsilon p^{2}_{w}(y)I&\;\;[P({\rm i}y)]^{\ast}\\[2.5pt] P({\rm i}y)&-\epsilon I\end{array}\right]&\quad\text{if }j=0\;,\\[13.00005pt] \left[\begin{array}[]{cc}0&\;-{\rm i}[P^{\prime}({\rm i}y)]^{\ast}\\[2.5pt] -{\rm i}P^{\prime}({\rm i}y)&\;\;0\end{array}\right]&\quad\text{if }j=1\;,\\[16.00008pt] \left[\begin{array}[]{cc}\epsilon(2w_{m}^{2}y^{2}+w_{c}^{2})I&\;-M^{\ast}\\[2.5pt] -M&\;\;0\end{array}\right]&\quad\text{if }j=2\;,\\[18.99995pt] 0&\quad\text{if }j=3\;,\\[13.00005pt] \left[\begin{array}[]{cc}-\epsilon w_{m}^{2}I&\;0\\ 0&\;0\end{array}\right]&\quad\text{if }j=4\end{array}\right.$
@@
1:  $\underline{z}\leftarrow$ a rightmost eigenvalue of $P(\lambda)$.
2:  $\underline{x}^{(0)}\leftarrow$ $\text{Re}(\underline{z})$.
@@
4:     Vertical Scan. Determine the intervals $[y^{L}_{j},y^{R}_{j}]$ for $j=1,\dots,l$ such that
@@
The sequence $\{x^{(k)}\}$ generated by Algorithm 1 is monotonically non-decreasing and satisfies $\>\lim_{k\rightarrow\infty}x^{(k)}=\alpha_{\epsilon}(P)$.
@@ 2.2 Minimization of the $\epsilon$-pseudospectral abscissa for small-scale problems
For simplicity, we combine the primal and dual variable into a vector $y=(s,\mu)$. Moreover, we denote by ${\mathcal{L}}_{s_{1}}$, ${\mathcal{L}}_{s_{2}}$ and ${\mathcal{L}}_{y}$, ${\mathcal{L}}_{yy}$ the partial derivative of ${\mathcal{L}}$ with respect to $s_{1}$, $s_{2}$ and the gradient, the Hessian of ${\mathcal{L}}$ with respect to $y$. If the optimizer $s(\nu)$ of (13) is unique and $\sigma(\nu,s(\nu))$ is simple, then, assuming the satisfaction of a constraint qualification, there exists a unique $\mu(\nu)$ such that ${\mathcal{L}}_{y}(\nu,s(\nu),\mu(\nu))=0$; see the proof of Theorem 2.6 below. We also combine the optimal values of the primal and dual variables, and write $y(\nu)=(s(\nu),\mu(\nu))$.
@@
for $k=1,\dots,\ell$, where $u_{j},v_{j}$ denote consistent unit left, respectively right singular vectors of $A(\widetilde{\mu})$ corresponding to the singular value $\sigma_{j}(\widetilde{\mu})$.
@@
Also, $g_{s_{1}}$, $g_{s_{2}}$ represent the partial derivatives of $g$ with respect to $s_{1}$, $s_{2}$, and $g_{s}$ the gradient of $g$ with respect to $s$.
@@
and $u,v$ are consistent unit left, right singular vectors, respectively, of $P(\widetilde{\nu},s(\widetilde{\nu}))$ corresponding to its smallest singular value $\sigma(\widetilde{\nu},s(\widetilde{\nu}))$.
@@
As $\widetilde{\nu}$ is non-degenerate, (13) with $\nu=\widetilde{\nu}$ has a unique optimizer, which we denote by $s(\widetilde{\nu})$, and $\sigma(\widetilde{\nu},s(\widetilde{\nu}))$ is simple. If $\sigma(\widetilde{\nu},s(\widetilde{\nu}))=0$ holds, then, by the continuity of $\sigma(\widetilde{\nu},s)$ with respect to $s$, there is a ball :=${\mathcal{B}}:=\{s\in{\mathbb{R}}^{2}\;|\;\|s-s(\widetilde{\nu})\|\leq r\}$ with a positive radius $r$ centered at $s(\widetilde{\nu})$ such that $\sigma(\widetilde{\nu},s)\leq\epsilon$ for all $s\in{\mathcal{B}}$, which contradicts the optimality of $s(\widetilde{\nu})$ for (13) with $\nu=\widetilde{\nu}$. Consequently, $\sigma(\widetilde{\nu},s(\widetilde{\nu}))$ is simple and nonzero. By continuity, there is an open neighborhood $\widetilde{U}\subseteq{\mathbb{R}}^{\mathsf{d}}$ of $\widetilde{\nu}$ and an open neighborhood $\widehat{U}\subseteq{\mathbb{R}}^{2}$ of $s(\widetilde{\nu})$ such that the singular value function $\sigma(\nu,s)$ remains simple and positive in $\widetilde{U}\times\widehat{U}$, while $p_{w}(\|s\|)>0$ in $\widehat{U}$, so $\sigma(\nu,s)$ as well as $g(s;\nu)$ defined as in (15) are real analytic in $\widetilde{U}\times\widehat{U}$ by Lemma 2.5. We assume, without loss of generality, that $g_{s}(s;\nu)\neq 0$, indeed the existence of a positive real number $\gamma$ such that $\|g_{s}(s;\nu)\|\geq\gamma$, for all $(s,\nu)\in\widetilde{U}\times\widehat{U}$ (if necessary by choosing $\widetilde{U}$, $\widehat{U}$ smaller).
@@
It is evident from (18) that the Lagrange multiplier $\mu(\vartheta)$ corresponding to the maximizer $s(\vartheta)$ is unique. Letting, for $\nu\in U$, $y(\nu):=(s(\nu),\mu(\nu))$, indeed, the first-order necessary conditions imply ${\mathcal{L}}_{y}(\nu,y(\nu))=0$ for every $\nu\in U$. Note also that $s(\widetilde{\nu})$ must be the unique maximizer of (13) with $\nu=\widetilde{\nu}$.
On the other hand, since ${\mathcal{L}}_{yy}(\widetilde{\nu},y(\widetilde{\nu}))$ is invertible by the non-degeneracy of $\widetilde{\nu}$, the analytic implicit function theorem implies the existence of a neighborhood $U_{0}\subseteq U$ of $\widetilde{\nu}$ and a unique function $w:U_{0}\rightarrow\widehat{U}$ satisfying $w(\widetilde{\nu})=y(\widetilde{\nu})$ and ${\mathcal{L}}_{y}(\nu,w(\nu))=0$ for all $\nu\in U_{0}$. Moreover, the function $w$ is real analytic. By the uniqueness of $w$, the function $\nu\in U_{0}\mapsto y(\nu)$ is unique and $y(\nu)=w(\nu)$ for every $\nu\in U_{0}$. This also shows the real analyticity of $y(\nu)$ at every $\nu\in U_{0}$. In particular, the map $\nu\in U\mapsto\alpha_{\epsilon}(\nu)=s_{1}(\nu)$ is real analytic at $\nu=\widetilde{\nu}$.
@@
$\begin{split}&\sigma_{s_{1}}(\widetilde{\nu},s(\widetilde{\nu}))\;=\;\mathrm{Re}\left\{u^{\ast}\frac{\partial P}{\partial s_{1}}(\widetilde{\nu},s(\widetilde{\nu}))v\right\}\;=\;\mathrm{Re}\{u^{\ast}P^{\prime}(s(\widetilde{\nu});\widetilde{\nu})v\}\,,\\[1.99997pt] &\frac{\partial}{\partial s_{1}}\left[p_{w}(\|s\|)\right]\bigg{|}_{s=s(\widetilde{\nu})}\;=\;\frac{s_{1}(\widetilde{\nu})}{\|s(\widetilde{\nu})\|}\,p^{\prime}_{w}(\|s(\widetilde{\nu})\|)\;,\end{split}$
where, in the first line, we use the analytic formula in (14), in particular $u$, $v$ denote a pair of consistent unit left, right singular vectors corresponding to the singular value $\sigma(\widetilde{\nu},s(\widetilde{\nu}))$. It then follows from (18) that
@@ 3 Computation of the $\epsilon$-pseudospectral abscissa for large-scale problems
1:  $z^{(0)}\leftarrow$ an eigenvalue of $P(\lambda)$.
2:  $V_{0}\leftarrow$ a unit norm right singular vector corresponding to $\sigma_{\min}(P(z^{(0)}))$.
@@
Interpolation. Suppose that the computed rightmost point $z^{(\ell)}$ after the execution of line 4 at the $\ell$th iteration, $\ell\geq 1$, is such that $\sigma_{\min}(P(z^{(\ell)}))$ is simple and nonzero. Using the variational characterization of the smallest singular value [9], as well as the analytical formulas for the derivatives of singular value functions (in particular Lemma 2.5), it can be shown for Algorithm 2 that the Hermite interpolation properties
@@ 4 Large-Scale minimization of the $\epsilon$-pseudospectral abscissa
for a matrix $V\in{\mathbb{C}}^{\mathsf{n}\times r}$ whose columns form an orthonormal basis for a small $r\ll\mathsf{n}$ dimensional restriction subspace ${\mathcal{V}}$. We form $M_{i}V,\,C_{j}V,\,K_{\ell}V\in{\mathbb{C}}^{\mathsf{n}\times r}$ for $i=1,\dots,\mathsf{m}$, $j=1,\dots,\mathsf{c}$, $\>\ell=1,\dots,\mathsf{k}\,$ in advance, then minimize
@@
1:  Initialize. $\nu^{(0)}_{1},\dots\nu^{(0)}_{\eta}\leftarrow$ initially chosen points in $\underline{\Omega}$. $z^{(0)}_{j}\leftarrow$ a rightmost point in $\Lambda_{\epsilon}(P(\,\cdot\,;\nu^{(0)}_{j}))$ using Algorithm 2 for $j=1,\dots,\eta$. $v^{(0)}_{j}\leftarrow$ a right singular vector corresponding to $\sigma_{\min}(P(z^{(0)}_{j};\nu^{(0)}_{j}))$ for $j=1,\dots,\eta$. $V_{0}\leftarrow$ a matrix whose columns form an orthonormal basis for span$\{v^{(0)}_{1},\dots v^{(0)}_{\eta}\}$.
@@
at every iteration. As argued in Section 3, it is reasonable to assume both (ii) and (iii) have ${\mathcal{O}}(\mathsf{n}^{2})$ computational cost. However, Algorithm 2 itself computes the smallest singular value a few times, so (ii) is usually more expensive than (iii). Thus, the overall cost of Algorithm 3 is mainly determined by the cost of (ii) and (iii) times the number of iterations of Algorithm 3. For large $\mathsf{n}$, we expect (ii) to dominate, as (i) involves only small reduced problems, that is the overall computational cost may largely be determined by the number of calls to Algorithm 2.
@@ 4.1 Solution of reduced minimization problems
${\mathcal{L}}^{\mathcal{V}}(\nu,s,\mu)\;:=\;s_{1}-\mu\{\sigma^{\mathcal{V}}(\nu,s)-\epsilon p_{w}(\|s\|)\}\;.$
@@
As in the non-reduced case, a point $\nu\in\Omega$ is called a non-degenerate point of the restriction of $P(\,\cdot\,;\,\cdot\,)$ to a subspace ${\mathcal{V}}$ if there is a unique optimizer of (22), denoted as $s^{\mathcal{V}}(\nu)=(s^{\mathcal{V}}_{1}(\nu),s^{\mathcal{V}}_{2}(\nu))$, the singular value $\sigma^{\mathcal{V}}(\nu,s^{\mathcal{V}}(\nu))$ is simple, and the Hessian $\nabla^{2}_{yy}{\mathcal{L}}^{\mathcal{V}}(\nu,y^{\mathcal{V}}(\nu))$ is invertible.
@@
in particular in terms of its gradient with respect to $s$, which we denote by $g^{\mathcal{V}}_{s}$.
@@
and $u,v$ are consistent unit left, right singular vectors, respectively, corresponding to the smallest singular value $\sigma^{\mathcal{V}}(\widetilde{\nu},s^{\mathcal{V}}(\widetilde{\nu}))$ of $P^{V}(\widetilde{\nu},s^{\mathcal{V}}(\widetilde{\nu}))$.

### v3

@@ 1 Introduction
[…]
for given (parameter-independent) matrices $M_{i},C_{j},K_{\ell}\in{\mathbb{C}}^{\mathsf{n}\times\mathsf{n}}$, and real-analytic functions $\psi_{i},\zeta_{j},\kappa_{\ell}:\Omega\rightarrow{\mathbb{R}}$ for $i=1,\dots,\mathsf{m}$, $j=1,\dots,\mathsf{c}$, $\ell=1,\dots,\mathsf{k}$. In a context where the dependence of $P(\lambda;\nu)$ on $\nu$ is not important, we may leave out this explicit dependence on $\nu$, and simply write
@@ 2.1 Computation of the $\epsilon$-pseudospectral abscissa for small-scale problems
[…]
@@
[…]
@@
[…]
@@
4:  Vertical Scan. Determine the intervals $y^{L}_{j},y^{R}_{j}]$ for $j=1,\dots,l$ such that
@@
The sequence $\{x^{(k)}\}$ generated by Algorithm 1 is monotonically non-decreasing and satisfies $\lim_{k\rightarrow\infty}x^{(k)}=\alpha_{\epsilon}(P)$.
@@ 2.2 Minimization of the $\epsilon$-pseudospectral abscissa for small-scale problems
For simplicity, we combine the primal and dual variable into a vector $y=(s,\mu)$. Moreover, we denote by ${\mathcal{L}}_{s_{1}}$, ${\mathcal{L}}_{s_{2}}$ and ${\mathcal{L}}_{y}$, ${\mathcal{L}}_{yy}$ the partial derivative of $\mathcal{L}$} with respect to $s_{1}$, $s_{2}$ and the gradient, the Hessian of $\mathcal{L}$} with respect to $y$. If the optimizer (i.e. global maximizer) $s(\nu)$ of (13) is unique and $\sigma(\nu,s(\nu))$ is simple, then, assuming the satisfaction of a constraint qualification, there exists a unique $\mu(\nu)$ such that ${\mathcal{L}}_{y}(\nu,s(\nu),\mu(\nu))=0$; see the proof of Theorem 2.6 below. We also combine the optimal values of the primal and dual variables, and write $y(\nu)=(s(\nu),\mu(\nu))$.
@@
[…]
@@
Also, $g_{s_{1}}$, $g_{s_{2}}$ represent the partial derivatives of $g$ with respect to $s_{1}$, $s_{2}$, and $g_{s}$, $g_{ss}$ the gradient of $g$, respectively the Hessian of $g$ with respect to $s$.
Before presenting the result, we remark on the third condition for the non-degeneracy of $\nu$, which is in terms of the invertibility of the Hessian ${\mathcal{L}}_{yy}(\nu,y(\nu))$. Using the definition of ${\mathcal{L}}(\nu,y)$, it is immediate that
${\mathcal{L}}_{yy}(\nu,y(\nu))=\left[\begin{array}[]{cc}-\mu(\nu)g_{ss}(s(\nu);\nu)&-g_{s}(s(\nu);\nu)\\ -g_{s}(s(\nu);\nu)^{T}&0\end{array}\right].$
Assuming that the first two conditions for the non-degeneracy of $\nu$ hold and that $g_{s}(s(\nu);\nu)\neq 0$, we must have $\mu(\nu)\neq 0$ (as apparent from the proof of Theorem 2.6 below, more specifically from (18)). Then, if $g_{ss}(s(\nu);\nu)$ is invertible, the Hessian ${\mathcal{L}}_{yy}(\nu,y(\nu))$ must be invertible, due to the fact that
$\left[\begin{array}[]{cc}-\mu(\nu)g_{ss}(s(\nu);\nu)&-g_{s}(s(\nu);\nu)\\ 0&(1/\mu(\nu))g_{s}(s(\nu);\nu)^{T}g_{ss}(s(\nu);\nu)^{-1}g_{s}(s(\nu);\nu)\end{array}\right]$
is invertible, recalling that $g_{s}(s(\nu);\nu)\neq 0$, $\mu(\nu)\neq 0$. Hence, the first two conditions for the non-degeneracy of $\nu$ combined with the assumptions (i) $g_{ss}(s(\nu);\nu)$ is invertible, (ii) $g_{s}(s(\nu);\nu)\neq 0$ imply the invertibility of ${\mathcal{L}}_{yy}(\nu,y(\nu))$.
@@
and $u,v$ are consistent unit left, right singular vectors, respectively, of $)P(\widetilde{\nu},s(\widetilde{\nu}$)) corresponding to its smallest singular value $)\sigma(\widetilde{\nu},s(\widetilde{\nu}$)).
@@
[…]
@@
(18) ${\mathcal{L}}_{s_{1}}(\vartheta,s(\vartheta),\mu(\vartheta))\;=\;1-\mu(\vartheta)\left\{\sigma_{s_{1}}(\vartheta,s(\vartheta))-\epsilon\frac{\partial}{\partial s_{1}}\left[p_{w}(\|s\|)\right]\bigg{|}_{s=s(\vartheta)}\right\}=\;0\,.$
@@
[…]
@@
[…]
where, in the first line, we use the analytic formula in (14), in particular $u$, $v$ denote a pair of consistent unit left, right singular vectors corresponding to the singular value $)\sigma(\widetilde{\nu},s(\widetilde{\nu}$)). It then follows from (18) that
@@ 3 Computation of the $\epsilon$-pseudospectral abscissa for large-scale problems
[…]
@@
[…]
@@ 4 Large-Scale minimization of the $\epsilon$-pseudospectral abscissa
for a matrix $V\in{\mathbb{C}}^{\mathsf{n}\times r}$ whose columns form an orthonormal basis for a small $r\ll\mathsf{n}$ dimensional restriction subspace $\mathcal{V}$}. We form $M_{i}V,\,C_{j}V,\,K_{\ell}V\in{\mathbb{C}}^{\mathsf{n}\times r}$ for $i=1,\dots,\mathsf{m}$, $j=1,\dots,\mathsf{c}$, $\ell=1,\dots,\mathsf{k}\,$ in advance, then minimize
@@
[…]
@@
at every iteration. As argued in Section 3, it is reasonable to assume both (ii) and (iii) have ${\mathcal{O}}(\mathsf{n}^{2})$ computational cost. However, Algorithm 2 itself computes the smallest singular value a few times, so (ii) is usually more expensive than (iii). Thus, the overall cost of Algorithm 3 is mainly determined by the cost of (i) and (ii) times the number of iterations of Algorithm 3. For large $\mathsf{n}$, we expect (ii) to dominate, as (i) involves only small reduced problems, that is the overall computational cost may largely be determined by the number of calls to Algorithm 2.
@@ 4.1 Solution of reduced minimization problems
${\mathcal{L}}^{\mathcal{V}}(\nu,s,\mu)\;:=\;s_{1}-\mu\{\sigma^{\mathcal{V}}(\nu,s)-\epsilon p_{w}(\|s\|)\}\,$,
@@
As in the non-reduced case, a point $\nu\in\Omega$ is called a non-degenerate point of the restriction of $P(\,\cdot\,;\,\cdot\,)$ to a subspace $\mathcal{V}$} if there is a unique optimizer of (22), denoted as $s^{\mathcal{V}}(\nu)=(s^{\mathcal{V}}_{1}(\nu),s^{\mathcal{V}}_{2}(\nu))$, the singular value $\sigma^{\mathcal{V}}(\nu,s^{\mathcal{V}}(\nu))$ is simple, and the Hessian ${\mathcal{L}}^{\mathcal{V}}_{yy}(\nu,y^{\mathcal{V}}(\nu))$ is invertible.
@@
in particular in terms of its gradient with respect to $s$, which we denote by $g^{\mathcal{V}}_{s}$. We also use $g^{\mathcal{V}}_{ss}$ to denote the Hessian of $g^{\mathcal{V}}$ with respect to $s$. Note that, as in the non-reduced case, the first two conditions for the non-degeneracy of a point $\nu$ of the restriction of $P(\,\cdot\,;\,\cdot\,)$ to $\mathcal{V}$} combined with the assumptions that $g^{\mathcal{V}}_{ss}(s^{\mathcal{V}}(\nu);\nu)$ is invertible and $g^{\mathcal{V}}_{s}(s^{\mathcal{V}}(\nu);\nu)\neq 0$ imply the invertibility of ${\mathcal{L}}^{\mathcal{V}}_{yy}(\nu,y^{\mathcal{V}}(\nu))$.
@@
and $u,v$ are consistent unit left, right singular vectors, respectively, corresponding to the smallest singular value $)\sigma^{\mathcal{V}}(\widetilde{\nu},s^{\mathcal{V}}(\widetilde{\nu}$)) of $)P^{V}(\widetilde{\nu},s^{\mathcal{V}}(\widetilde{\nu}$)).